OpenStudy (anonymous):

Use guess-and-check to find two numbers that have the sum of 10 and a difference of 6

OpenStudy (anonymous):

x + y = 10 x - y = 6 ------------ 2x = 16 x=8 8 + y = 10 y = 2

OpenStudy (anonymous):

u were supposed to guess first @surdawi and then use the method that u did for only checking ur answer as the question says